OpenStudy (anonymous):

Multiply, and simplify if possible Sqrt(7x) { sqrt(x)-7 sqrt(7) }

OpenStudy (kuoministers):

Are you still there?

OpenStudy (anonymous):

yes Hallo

OpenStudy (anonymous):

I'll be here until I get help xD

OpenStudy (kuoministers):

\[Sqrt(7x) { \sqrt(x)-7 \sqrt(7) }\] \[{ x \sqrt(7)- 49 \sqrt(x) }\]

OpenStudy (kuoministers):

sqrt something times sqrt something is sqrt (two things multiplying each other)

OpenStudy (anonymous):

Thank you, thank you. You'r saving Algebraic lives!

OpenStudy (kuoministers):

Thanks!